Please, could you tell me, what are the molecular electron configurations of Methane, Methanol and Chloromethane?


1
Expert's answer
2021-09-23T04:51:44-0400

(1)Methane


"CH_4=C(6)+4\times H(1)\\=10 e^-\\"

so the electron configurations of Methane is

"\sigma1s^2\sigma^*1s^2 \sigma2s^2\sigma^*2s^2[\pi2px^1=\pi2py^1]"

(2) Methanol


"CH_3OH=C(6)+O(8)+4\times H(1)\\=15e^-"

so the electron configurations of Methanol is

"\sigma1s^2\sigma^*1s^2 \sigma2s^2\sigma^*2s^2 \sigma2pz^2[\pi2px^2=\pi2py^2][\pi^*2px^1]"

(3) Chloromethane


"CH_3CL=C(6)+3\times H(1)+CL(9)\\=16e^-"

so the electron configurations of Chloromethane is

"\sigma1s^2\sigma^*1s^2 \sigma2s^2\sigma^*2s^2 \sigma2pz^2[\pi2px^2=\pi2py^2][\pi^*2px^1=\pi^*2py^1]"